Outer electrons Inner core electrons Element Valence electrons Fe Cu Cre Bi

list the Inner core electrons, the outer electrons and the valence electrons.
Answer #1

1) F - atomic number is 9

Configuration of F - 1s2, 2s2, 2p5

Inner core electrons = 2

Outer core electrons = 5

Valence electrons corresponding maximum principal quantum number i.e. 2 here =  2+5 = 7

2) Atomic number of Cu is 29

Configuration - 1s2, 2s2, 2p6, 3s2, 3p6 , 4s1, 3d10

Inner core - 18 electrons

Valence electrons - 1

Outer core - 10

3)

Atomic number of Cr is 24.

Configuration - 1s2, 2s2, 2p6, 3s2, 3p6, 4s1, 3d5

Valence electron = 1

Inner core electrons - 18

Outer core electrons - 5

4) Atomic number of Bi is 83

Configuration - 1s2, 2s2, 2p6, 3s2, 3p6, 4s2, 3d10, 4p6, 5s2, 4d10, 5p6, 6s2, 4f14, 5d10, 6p3

Valence electron corresponding to 6 = 2+3 = 5

Inner core electrons = 78

Outer core is 6p and having electrons = 3

Let me know if any doubts.